OpenStudy (somersetday):

Dora drew JKL and MNP so that K N, JK = 6, MN = 18, KL = 4, and NP = 12. Are JKL and MNP similar? If so, identify the similarity postulate or theorem that applies. A. Cannot be determined B. Similar - SAS C. Similar - AA D. Similar - SSS

OpenStudy (aakashtomar):

This is a really nice and simple problem on similarity. You just need to know about the basic requirements to call a similarity as SAS, AAA etc. Okay, if this is all what is given, ( assuming you've given all the info given in the question ) it cannot be determined. It could've been similar had it been given that angle K = angle N. If that is true, we can easily use SAS similarity here. Observe that JK:MN = 1/3. Also KL:NP=1/3. And if angle K = angle N then it is SAS as it is the angle between the sides which are in the same ratio. :)